-) f(n)=n+1 g(n) = 3n² – 3 Find (fºg)(8)
Answer #1

) f(n)=n+1 g(n) = 3n² – 3 Find (fºg)(8) · fog (n) = 3n-3 +1 tog(n) = 3nt-2 fog (8) = 3.8?- 2 = 190